An inert gas in a closed system initially at A undergoes changes in pressure (P) and volume (V) and returns back to A, ac
Answer #1

The energy change per stage is calculated:

A-B:

E = P * dV + V * dP = 10 * (4-1) + 1 * (40-10) = 60 Pa * m3

B-C:

E = P * dV = 40 * (1-4) = - 120 Pa * m3

AC:

E = V * dP = 1 * (10-40) = - 30 Pa * m3

The total energy change:

Etotal = 60 - 120 - 30 = - 90 J * (1 cal / 4.18 J) = - 21.53 cal

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.